OpenStudy (ryanchick):

i will fan and medal 4) (5)(3-1)+4

pooja195 (pooja195):

is 4 part of the problem?

OpenStudy (raffle_snaffle):

is that 4) representing the 4th problem or is that part of the question?

OpenStudy (ryanchick):

that the whole problem

pooja195 (pooja195):

so (4) (5)(3-1)+4 correct?

OpenStudy (ryanchick):

ya now i just need to figure the problem out

pooja195 (pooja195):

Use PEMDAS P is first (parentheses) so 3-1=?

OpenStudy (ryanchick):

2

pooja195 (pooja195):

Good now we have (4)(5)(2)+4 5 x 4=?

OpenStudy (ryanchick):

20

pooja195 (pooja195):

Good now we have 20(2)+4 20 x 2=?

OpenStudy (ryanchick):

40

pooja195 (pooja195):

40+4=?

OpenStudy (ryanchick):

44

pooja195 (pooja195):

good that's our final answer

OpenStudy (ryanchick):

thxs

pooja195 (pooja195):

yw